In a government expenditure multiplier, if autonomous spending a is 1000, the marginal propensity to save is 0.25 and the marginal tax rate t is 0.30. what will be the equilibrium level of income?


1
Expert's answer
2020-08-18T12:49:07-0400

Marginal propensity to consume =1-0.25=0.75

Tax=0.3"\times"1000=300


Equilibrium level of income is reached when Y=AE

AE=C+I =1000+0.75(Y-300)

=1000+0.75Y-225

Setting the equilibrium ;

AE=775+0.75Y


0.25Y=775

Y="\frac{775}{0.25}"


Y=3,100